OpenStudy (howard-wolowitz):

Urgent! one question :0

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{x+\frac{4y}{x} }{\frac{7}{3x}} }\) this is how I would start, \(\large\color{black}{ \displaystyle \frac{\frac{x^2}{x}+\frac{4y}{x} }{\frac{7}{3x}} }\)

OpenStudy (michele_laino):

please what is: \[x + \frac{{4x}}{y} = ...?\]

OpenStudy (solomonzelman):

oh, my bad

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{x +\frac{4x}{y}}{\frac{7}{3x}} }\)

OpenStudy (solomonzelman):

combine the fractions, apply the rule that